# Accelerating Parallel Diffusion Model Serving with Residual Compression

### v1

@@ 4.3 Ablation Studies
Like displaced parallel, CompactFusion requires at least a 1-step warmup, where the uncompressed activation is used to initialize the base tensor for later residual computation (detailed in Appendix C). Figure 8 compares different methods and warmups. We observe that CompactFusion maintains stable and high visual quality with just a single warmup step, showing little degradation compared to longer warmup. In contrast, baseline methods are more sensitive to warmup configuration. For more detailed results, please refer to Section E.3
@@
Error Feedback significantly improves quality. For Compact-1bit under 1-step warmup, it reduces FID from 19.23 to 7.08, lowers LPIPS from 0.389 to 0.260, and increases PSNR from 19.78 to 22.90. For more detailed results, please refer to Section E.2.
@@ Appendix A Limitations
Appendix B Discussion
@@
Appendix C Implementation Details
@@
Similar to displaced parallelism, CompactFusion requires at least one warmup step. It is used to initialize the uncompressed base tensor for residual calculation. Ablation studies show that CompactFusion is highly robust and achieves strong generation quality even with a single warmup step (Section 4.3). There are two practical ways to implement warmup: (1) running the warmup steps with parallelism but without compression; or (2) letting each device execute the warmup locally without any communication. The first method introduces some communication but is typically faster on PCIe and NVLink. However, it may become a bottleneck under extremely slow networks. The second method avoids communication but incurs slightly higher latency due to lack of parallelism. We adopt the first method in our experiments.
@@
To compute this approximation efficiently, we use a -step subspace iteration [26]; subspace iteration was first used in gradient compression [27]:
Appendix D More Experiment Details
D.1 Human Evaluation
@@
Each question displayed one reference image (generated by the original uncompressed model) and five test images produced by different parallel strategies (including DistriFusion, PipeFusion, and CompactFusion variants), as shown in Figure 9. The five options were randomly shuffled to avoid position bias. The prompts used for generation were randomly sampled from the COCO dataset.
@@ D.2 More Setups
The default scheduler used is DPM solver. [37]
@@
Appendix E More Quantitative Results
@@
E.1 Effect of Tradeoff between Rank and Precision
@@
Compact-Lowrank uses rank 32 and applies INT4 quantization to the transmitted and , matching the communication cost of rank-8 FP16 methods. In contrast, the Pure Lowrank baselines use lower ranks (8 or 12) and do not quantize. As shown in Table 3, increasing the subspace iteration steps (from 2 to 10) brings negligible quality gains, while increasing the rank significantly improves generation quality. This highlights the importance of directional coverage over projection precision under tight bandwidth budgets.
E.2 Effect of the Error Feedback
@@
The results (Table 4) show that incorporating error feedback significantly improves performance across all metrics. Specifically, the Compact-1bit with Error Feedback variant achieves higher PSNR scores, lower LPIPS values, and notably better FID scores compared to the version without error feedback. This indicates that error feedback helps preserve perceptual and structural quality, making it a critical component for effective low-bit communication.
E.3 Effect of the Warmup steps
@@
Table 5 demonstrates that longer warm-up phases and higher-bit compression with error feedback improve visual quality.
E.4 Details on Human Evaluation Results
@@
Compact-2bit method can exhibit high consistency and fidelity with regards to human perception (Table 6).
Appendix F More Qualitative Results
@@
F.1 FLUX.1-dev Qualitative Results
The supplementary qualitative results are shown in Table 7.
@@
F.2 CogVideoX-2b Qualitative Results
@@
Appendix G Compression Error Analysis
G.1 Analysis Overview
@@ G.3 Assumptions
Assumption 1
Appendix H License and Asset Attribution
@@
Appendix I Acknowledgments
We thank the xDiT team for providing a comprehensive parallel diffusion framework, which made this work possible.

### v2

@@ 4.3 Ablation Studies
Like displaced parallel, CompactFusion requires at least a 1-step warmup, where the uncompressed activation is used to initialize the base tensor for later residual computation (detailed in Appendix˜D). Figure˜8 compares different methods and warmups. We observe that CompactFusion maintains stable and high visual quality with just a single warmup step, showing little degradation compared to longer warmup. In contrast, baseline methods are more sensitive to warmup configuration. For more detailed results, please refer to Section˜F.3
@@
Error Feedback significantly improves quality. For Compact-1bit under 1-step warmup, it reduces FID from 19.23 to 7.08, lowers LPIPS from 0.389 to 0.260, and increases PSNR from 19.78 to 22.90. For more detailed results, please refer to Section˜F.2.
@@ 5 Conclusion
6 Acknowledgements
This work is supported in part by the National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204 and 62472249), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001). This study is also supported in part by the Natural Science Foundation of Top Talent of SZTU(Grant No. GDRC202413).
We would like to express our sincere appreciation to the xDiT team for their comprehensive parallel diffusion framework, which served as the foundation for our experiments.
@@ Appendix A Limitations
Appendix B More Preliminaries
B.1 Sequence Parallelism
Sequence parallel enables multiple GPUs to collaboratively generate a single image or video sample, thereby reducing the latency per sample. During inference, an image or video frame is first divided into a sequence of patches (tokens). Sequence parallel partitions this sequence of patches across multiple devices so that each GPU is responsible for computing a section of it.
After the partition, however, inter-GPU communications are required. For FFN layers, communication is unnecessary—each patch (token) is processed independently, and all required data resides locally on the same device. But for attention layers, every patch must attend to all other patches across the full sequence, which introduces cross-device dependencies.
Different parallel strategies handle this communication differently:
Patch Parallel[distrifusion]. This strategy performs a single AllGather operation over all devices to collect the complete set of keys and values (Figure˜9). Each GPU thus receives global and tensors before executing the attention computation. It may incur substantial communication overhead as the entire matrices are transmitted every step.
Ring Attention[ring]. Instead of a full AllGather, Ring Attention organizes the devices into a logical ring and performs a series of peer-to-peer exchanges. Each stage transmits only a shard of and computes the corresponding partial result. Computation and communication are interleaved, offering partial overlap.
Ulysses[ulysses]. Ulysses adopts a slightly different design, using two All-to-All operations to exchange the required data between devices.
We provide illustrative pseudocode in Algorithms˜1 and 2.
B.2 Displaced Patch Parallel (DistriFusion)
Patch Parallel performs a synchronous AllGather of at every denoising step and waits immediately before attention, which leads to substantial blocking when the transfer payload is large. Displaced Patch Parallel [distrifusion] avoids this by adopting a send-now, receive-later schedule with a one-step displacement. In each iteration, the system launches a non-blocking communication for the current but computes attention using the previous step’s gathered , thereby overlapping communication with computation. Each iteration waits on the handle from the earlier launch when those tensors are actually required. This schedule reduces latency at the cost of using stale keys and values, which can introduce quality drift.
Appendix C Discussion
@@
Appendix D Implementation Details
@@
Similar to displaced parallelism, CompactFusion requires at least one warmup step. It is used to initialize the uncompressed base tensor for residual calculation. Ablation studies show that CompactFusion is highly robust and achieves strong generation quality even with a single warmup step (Section˜4.3). There are two practical ways to implement warmup: (1) running the warmup steps with parallelism but without compression; or (2) letting each device execute the warmup locally without any communication. The first method introduces some communication but is typically faster on PCIe and NVLink. However, it may become a bottleneck under extremely slow networks. The second method avoids communication but incurs slightly higher latency due to lack of parallelism. We adopt the first method in our experiments.
@@
To compute this approximation efficiently, we use a -step subspace iteration [subspace]; subspace iteration was first used in gradient compression [powersgd]:
Residual Base and Error Feedback.
In CompactFusion, each transmitted tensor represents a residual, which is the change in activation relative to a stored base. The base is initialized with the uncompressed activation during the warm-up step and updated after every iteration. Compression introduces an error that is observable only on the sender side, since the receiver never sees the true residual before compression. Therefore, error feedback is applied at the sender: the previously accumulated error is added back to the next residual before compression, ensuring that lost information is gradually compensated over time.
Integration into Parallel Pipelines.
CompactFusion integrates into existing parallel frameworks by simply wrapping standard communication primitives such as send/recv and AllGather. This modular design allows the same compression interface to be reused across different parallel strategies, ensuring compatibility with existing diffusion inference pipelines.
Appendix E More Experiment Details
E.1 Human Evaluation
@@
Each question displayed one reference image (generated by the original uncompressed model) and five test images produced by different parallel strategies (including DistriFusion, PipeFusion, and CompactFusion variants), as shown in Figure˜10. The five options were randomly shuffled to avoid position bias. The prompts used for generation were randomly sampled from the COCO dataset.
@@ E.2 More Setups
The default scheduler used is DPM solver. [dpmsolver]
@@
Appendix F More Quantitative Results
@@
F.1 Effect of Tradeoff between Rank and Precision
@@
Compact-Lowrank uses rank 32 and applies INT4 quantization to the transmitted and , matching the communication cost of rank-8 FP16 methods. In contrast, the Pure Lowrank baselines use lower ranks (8 or 12) and do not quantize. As shown in Table˜3, increasing the subspace iteration steps (from 2 to 10) brings negligible quality gains, while increasing the rank significantly improves generation quality. This highlights the importance of directional coverage over projection precision under tight bandwidth budgets.
F.2 Effect of the Error Feedback
@@
The results (Table˜4) show that incorporating error feedback significantly improves performance across all metrics. Specifically, the Compact-1bit with Error Feedback variant achieves higher PSNR scores, lower LPIPS values, and notably better FID scores compared to the version without error feedback. This indicates that error feedback helps preserve perceptual and structural quality, making it a critical component for effective low-bit communication.
F.3 Effect of the Warmup steps
@@
Table˜5 demonstrates that longer warm-up phases and higher-bit compression with error feedback improve visual quality.
F.4 Details on Human Evaluation Results
@@
Compact-2bit method can exhibit high consistency and fidelity with regard to human perception (Table˜6).
F.5 Temporal Stability Evaluation
Method Temporal Flickering (↑) Motion Smoothness (↑) Original (No Compression) 0.9721 0.9817 DistriFusion 0.9658 0.9795 Compact-2bit 0.9644 0.9808 Compact-1bit 0.9661 0.9806 Compact-Lowrank 0.9623 0.9775
The results (Table˜7) show that all CompactFusion variants maintain temporal stability comparable to or better than prior methods. The differences across metrics are minimal, indicating that compression introduces no noticeable degradation in temporal stability.
Appendix G More Qualitative Results
@@
G.1 FLUX.1-dev Qualitative Results
The supplementary qualitative results are shown in Table˜8.
@@
G.2 CogVideoX-2b Qualitative Results
@@
Appendix H Compression Error Analysis
H.1 Analysis Overview
@@ H.3 Assumptions
Assumption 1 (-Smoothness). The function is -smooth with
which implies that the process contracts over time, ensuring stability.
Assumption 2 (-Compressor). The compressor satisfies
where measures the compression quality (higher indicates better fidelity).
Assumption 3 (Bounded True Residual Variance). The true residuals have bounded expected energy:
where is a finite constant.
Assumption 4 (Bounded Activation Variance). The target activations have bounded expected values:
where is a finite constant.
Assumption 5 (Uncorrelated Errors (Simplification)). We assume certain terms are uncorrelated in expectation:
(true residual uncorrelated with previous total error),
(perturbation effect uncorrelated with true residual),
(perturbation effect uncorrelated with previous total error),
Compression error and propagation error are uncorrelated.
Here, is the perturbation effect due to the error in the input to
H.4 Analysis Model and Scope
Throughout this analysis, we model the evolution of the system state using the reconstructed state from the previous step, i.e., . This means the function always operates on the potentially noisy state resulting from previous compression steps.
In practical parallel implementations, such as sequence parallelism, a device might compute attention using a combination of locally available, uncompressed ("fresh") Key/Value pairs and received Key/Value pairs that have undergone compression/decompression. Our analysis, by assuming the entire input context has passed through the compression cycle, simplifies the setup and likely provides a conservative upper bound on the error compared to such mixed-context scenarios. The core principles of error propagation and the benefits of compressing smaller residuals, however, remain applicable.
H.5 Error Analysis for Naive Compression
Proposition H.1.
Let denote the steady-state mean squared error upper bound under naive compression. Assuming the process has reached stationarity, the error satisfies the bound
Proof.
In the naive scheme, . The total error is decomposed into compression error and propagation error . Assuming these errors are uncorrelated (section˜H.3):
Using the -compressor property (section˜H.3) and the bounded activation variance assumption (section˜H.3), we can bound the compression error by
Using the -smoothness assumption (section˜H.3), we can bound the propagation error by
Combining the terms yields the recurrence
Letting as the steady-state error upper bound and assuming in steady state, we obtain:
Solving for gives the bound
where stability requires .
H.6 Error Analysis for Residual Compression
Proposition H.2.
Let denote the steady-state mean squared error upper bound for residual compression. Assuming the process has reached stationarity, the error is bounded by
Proof.
In the residual scheme, we let , , and . The total error is , where . Follow the assumption that and are uncorrelated (section˜H.3), we can get
Using the -compressor property (section˜H.3):
Since
where and . Applying the uncorrelatedness assumptions (section˜H.3), we can further bound
Thus, the compression error is bounded by
The propagation error term is bounded exactly as in the naive case: .
Combining the bounds for compression and propagation errors:
Letting as the steady-state error upper bound and assuming in steady state, we obtain:
Solving yields the closed-form upper bound:
where stability requires:
H.7 Analysis Without Error Feedback
Consider a hypothetical scheme where the true residual is compressed and added to the previous reconstruction: . The error evolves as:
where is the compression error on the true residual. If we assume are uncorrelated over time, the expected squared error (variance) accumulates:
This indicates that grows linearly with and does not converge to a finite steady state, highlighting the necessity of the error feedback mechanism (implicit in using to compute the target and the residual ) for stability.
H.8 Steady-State Error Bounds
We present the derived steady-state expected squared error bounds for naive compression () and residual compression with error feedback (), based on the given assumptions. The results are as follows.
Provided the stability condition for residual compression is satisfied and given that, typically, , the residual scheme offers a lower theoretical error bound (although a precise comparison requires evaluating all expressions, including denominators).
Appendix I License and Asset Attribution
@@
xDiT(xfuser) [xfuser]: A parallel diffusion inference framework released under the Apache-2.0 license.
DistriFusion(distrifuser) [distrifuser]: Released under the MIT license.